OpenStudy (anonymous):

To solve the equation four r equals two-thirds, you should: HELP PLEASE!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (godlovesme):

4r=2/3 Is this ur question? @DivergentMockingjay19

OpenStudy (anonymous):

yes

OpenStudy (godlovesme):

The easy way yo solve for r is 1st reduce 2/3 --> .66667 Now we have 4r= .66667 can u solve from there? :(

OpenStudy (godlovesme):

:) ***

OpenStudy (anonymous):

ummm...I am confused?

OpenStudy (godlovesme):

4r =2/3 if u divide 2 by 3 u will get .66667 right? 4r= .66667 divide both side by 4 and r will be .1666 BTW do u have options?

OpenStudy (anonymous):

oh I am getting it now... :)

OpenStudy (godlovesme):

Great sorry for the confusion =0

OpenStudy (anonymous):

it fine